Question 5 (20 marks)

Part A (13 marks)

The Rockwood Co. just paid a dividend of 1.5 per share of stock. The following table lists the Rockwood Co.’s earnings per share and dividend per share over time.

a) Assume the company strictly follows the Lintner model to develop its dividend payout policy. Please fill in the following table and show your steps according to Lintner model (Calculate the dividend three/four/five year from now).

b) Is dividend changes lead or lag earnings change? Briefly explain it.

Time

0

1

2

3

4

5

EPS

4

4

4.8

4.8

4

4

Div

1.5

1.6

1.8

Div(3)?

Div(4)?

Div(5)?

Part B (7 marks)

You own 4,000 shares of stock in Burklin Corporation. You will receive a $10 per share cash dividend in one year. In two years, Burklin will pay a liquidating dividend of $90 per share. The required return on Burklin’s stock is 12 percent. Ignore taxes and transaction costs.

a) What is the current share price of your stock?   (2 marks)

b) If you would rather have equal dividends in each of the two years, show how you can accomplish this by creating homemade dividends. (5 marks)
Answer #1

Part A :

a.

Dividend paid for the current year = 1.5 per share

Here, given the EPS that is Earnings per Share and DPS that is, Dividend per share, we have to calculate the Dividend payout Ratio.

Dividend Payout Ratio = Dividend per share/Earnings per share

Year 0 = 1.5/4 = 0.375

Year 1 = 1.6/4 = 0.4

Year 2 = 1.8/4.8 = 0.375

Year 3 would be in the ratio of year 1 as dividends are issued in the ratio of alternative years,

Year 3 = 0.4 x 4.8 = 1.92

Year 4 = 0.375 x 4 = 1.5

Year 5 = 0.4 x 4 = 1.6

b.

Generally when dividends are distributed , they tend to lag behind the Earnings. That is, increase in dividend follows increase in earnings. Thus dividends tend to lag behind the earnings and hence this causes regressing earnings on lagged dividends. This further concludes that the dividends have to be distributed based on the firm specific characteristics as well.

PART B

a. current dividend price = 10

number of years = 2

current stock price = D1/(1+Ke)^n

D1 is the current dividend price

Ke is the time value of money

n is the number of years

Hence current stock price = ( 100/1+1.12)^2

Hence current stock price = $80 per share

B. Homemade dividends are those the an individual earns as a form of investment income by partially selling a part of his portfolio. Here hence first we have to reduce the current dividend from the stock price that is ,

$80 -$10 = $70

Thus now the investor will have 4000 shares @ $70 per share which will equate to $280000 which is the ex dividend price and hence it will divide it equally over the two years.